OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x -2 0 4 f(x) -7 3 -73 A. y = -4x2 - 3x + 3 B. y = 4x2 + 3x - 3 C. y = -3x2 - 3x + 4 D. y = -3x2 - 4x + 3

OpenStudy (anonymous):

Replace the values of x into the 4 equations to find the one that fits. Exp A: -4(-2)^2 - 3(-2) + 3 = -16 + 6 + 3 = -7. OK Check: x = 0 -> y = 3. OK x = 4 ->y = -64 - 12 + 3 = -73. OK

OpenStudy (anonymous):

i understand what you are doing on the first part (exp a) i dont understand hwta you did when you put check... sorry i am terrible at math :(

OpenStudy (anonymous):

what*

OpenStudy (anonymous):

I used the first column data (x = -2; y= -7) to test with equation A. The data fit, so A is the answer. I verified the data of other columns (x = 0, x = 4) to see if the equation A does also fit.

OpenStudy (anonymous):

holy crap i understand haha thank you

OpenStudy (anonymous):

how do i close the question and give you a medal?